OpenStudy (anonymous):

A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 2y ft. and the middle length side x + 5y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

SOMEONE HELP

OpenStudy (anonymous):

Are of a triangle is 1/2 base x perpendicular height. Take the base to be the shortest length. The hypotenuse is always the longest so the middle length is our perpendicular height. So we get: 1/2 X (x-2y) X (x+5y) = 1/2 X (x^2+3xy-10y^2) =\[\left( x^2+3xy-10y^2 \right)\div 2\]